OpenStudy (anonymous):

Find the exact value of cos135degrees. Tell the quadrant

OpenStudy (kohai):

Have you ever used a unit circle before?

OpenStudy (anonymous):

Yes

OpenStudy (kohai):

Ok, awesome. Pull one up. What is the coordinate at 135 degrees?

OpenStudy (anonymous):

3pi/4

OpenStudy (kohai):

That's the converted pi value. We need the actual coordinates

OpenStudy (kohai):

Do you know where the coordinates are?

OpenStudy (anonymous):

I dont have that on my chart yet

OpenStudy (kohai):

Okay. That's probably why you're having issues with this. Take a look at this and fill in your chart http://etc.usf.edu/clipart/43200/43215/unit-circle7_43215_lg.gif

OpenStudy (kohai):

So essentially, the coordinates are (cos, sin). So if at pi/3 degrees, the coordinates are (1/2, sqrt3/2), then cos=1/2 and sin=sqrt3/2. Does that make sense?

OpenStudy (anonymous):

I think I get it

OpenStudy (kohai):

So at the coordinate of 135, what is the value of cos?

OpenStudy (anonymous):

sqroot2/2?

OpenStudy (kohai):

Careful, remember the coordinates are (cos, sin)

OpenStudy (kohai):